OpenStudy (anonymous):

a bag has 5 red marbles, 6 blue marbles and 4 black marbles. What is the probability of picking a red marble, replacing it, and then picking a blue marble

OpenStudy (johnweldon1993):

Well if they marbles are replaced...the chances would not change from each turn... In the beginning...we have 5 red...out of 15 total marbles...so the probability would be 5/15 or 1/3 And the next time...the probability is 6 out of 15 (since we didnt take one away) so out probability is now 6/15 So our probability is \[\large \frac{1}{3} \times \frac{6}{15}\]